# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 5.3.2. There exists an old individual with large offspring
The final ingredient to prove persistence is to show that there exists an ‘old’ individual that survives until time $\approx t$ and has a sufficiently large offspring. Combined with the result in Lemma 5.9, this implies that the birth-time of the individual with the largest offspring (i.e. $I_{t}^{\mathrm{ct}}$) cannot be ‘too large’ compared to $O_{t}^{\mathrm{ct}}$. We make this precise in the following result.
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\right)\geq 1-\varepsilon.$ (5.60)
@@
[…]
@@
[…]
@@
[…]
×$\times\exp\big{(}-(d^{*}+\eta)(2A+\tfrac{1}{2})C_{1}\big{)}.$
@@
[…]
@@
$c_{\infty}(A,C_{1}):=\frac{1}{2}\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right),$ (5.67)
@@
[…]
Using Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$, we can bound the right-hand side from above for all $t>t^{\prime\prime}$ by
@@
$\mathbb{P}_{\mathcal{S}}(\forall{}$ $s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k)$ (5.74)
@@
[…]
@@
[…]
@@ 5.4. Persistence in the PAVD model
Fix $\varepsilon,A>0$, $t^{\prime}>0$ large (according to Proposition 5.8, Lemma 5.9, and Lemma 5.10) and recall $k=k(C_{1})$ and $u_{t}$ from (5.48). Then, for any $t>t^{\prime}$:
@@
    We take $K>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$: $O_{s}^{\mathrm{ct}}>u_{t}-K$.
@@
    We take $C_{1}>0$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)\geq k$.
@@
    We take $C_{2}>\tfrac{1}{2}C_{1}$ sufficiently large so that, with probability at least $31-\varepsilon/$ 3, for all $s\in[t-A,t+A]$ $\forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}$: $d_{v}(s)<k$.
$(a)$, $(b)$, and $(c)$ follow from Proposition 5.8, Lemma 5.10, and Lemma 5.9, respectively. Then, $(b)$ and $(c)$ combined imply that at $s$ the alive individual with the largest offspring has birth-time at most $u_{t}+C_{2}$, for all $s\in[t-A,t+A]$, with probability at least $1-\tfrac{2}{3}\varepsilon$. That is, $I_{s}^{\mathrm{ct}}<u_{t}+C_{2}$ for all $s\in[t-A,t+A]$ with probability at least $1-\tfrac{2}{3}\varepsilon$. Combined with $(a)$, we thus conclude that the result follows for $M=C_{2}+K$. ∎
We finally prove Theorem 2.6.
@@
We let $F_{n,A}:=\big{[}\frac{1}{\lambda^{*}}\log n-A,\frac{1}{\lambda^{*}}\log n+A\big{]}$ for ease of writing. Then,
@@
[…]
@@
The second term on the right-hand side is at most $\varepsilon/3$ for all $n\geq n_{0}$ by (5.78). Using Proposition 5.8 with $t=\frac{1}{\lambda^{*}}\log n$ we can take $M=M(A,\varepsilon)$ and $n_{1}=n_{1}(A,\varepsilon)\geq n_{0}$ such that the first term on the right-hand side is at least $31-\varepsilon/$ 3 for all $n\geq n_{1}$ as well, so that the entire lower bound is at least $1-\frac{2}{3}\varepsilon$. We then define
$\ell_{n}:=\frac{d^{*}}{\lambda^{*}+d^{*}}\log n+\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big{(}r\big{(}\tfrac{1}{\lambda^{*}}\log n\big{)}\big{)},$ (5.80)
@@
[…]
$-\mathbb{P}_{\mathcal{S}}\left(N\Big{(}\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\geq\mathrm{e}^{M^{\prime}+\ell_{n}}\right).$
The first probability is at least $1-\frac{2}{3}\varepsilon$ by taking $M^{\prime}=2\lambda^{*}M$ and applying (5.79) with $M$ large. Then, since $N(t)\geq|\mathcal{B}((0,t))|$ for all $t\geq 0$ deterministically,
$\mathbb{P}_{\mathcal{S}}\left(N\Big{(}-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(\Big{|}\mathcal{B}\Big{(}0,-\frac{M^{\prime}}{2\lambda^{*}}+\frac{1}{\lambda^{*}}\ell_{n}\Big{)}\Big{|}\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ (5.83)
@@
We continue by proving that $I_{n}/O_{n}$ is tight. We apply Proposition 5.11 to conclude that there exist $n_{1}\geq n_{0}$ and $M=M(\varepsilon,A,n_{1})>0$ such that for all $n\geq n_{1}$,
@@
We can bound the first term on the right-hand side from below by $31-2\varepsilon/$ 3 for all $n\geq n_{0}$ using (5.84) and setting $M^{\prime}:=C^{2}\mathrm{e}^{\lambda^{*}M}$. The second term can be bounded from above by using that $N(t)\leq 2|\mathcal{B}(0,t)|$ deterministically for all $t\geq 0$, so that
@@
[…]
@@
The final term is at most $\varepsilon/2$ for $A$ sufficiently large by (5.78). Then, the first probability on the right-hand side is also at least $21-\varepsilon/$ 2 by setting $C_{1}^{\prime}=K$ in Lemma 5.9 and $C_{1}=K$ in Lemma 5.10 and taking $K$ sufficiently large (with $t=\frac{1}{\lambda^{*}}\log n$ and $n$ sufficiently large). We thus arrive at the desired result and conclude the proof. ∎
@@ 6. Discussion and Open Problems
Existence of persistent $m$-hubs for more general CMJ branching processes. In Section 4 we studied the existence of persistent $m$-hubs in CMJ branching processes. Here, we were able to provide general conditions for the non-existence of persistent $m$-hubs, but could only prove the existence of persistent $1$-hubs (i.e. a persistent maximum degree) under general conditions. For persistent $m$-hubs with $m\geq 2$ we required the inter-birth times to be exponentially distributed, as the proofs leveraged the memoryless property at a few places. This leads to the following open problem.
@@
Can Theorem 4.4 be extended to hold for more general inter-birth random variables that satisfy Assumption 4.1?
Furthermore, the results in Theorems 4.2 and 4.4 are restricted to the setting in which
$\sum_{i=1}^{\infty}X(i)=\infty\text{ almost surely and }\widehat{\mu}(\lambda)<1\text{ for some $\lambda>0$}.$ (6.1)
The existence of persistent $1$-hubs in CMJ branching processes that do not satisfy at least one conditions does exist, see e.g. [33, 26, 25, 31], but these works restrict themselves to the case $D=\infty$ almost surely. Can such results be proved in more general settings when $\mathbb{P}\!\left(D=\infty\right)\in[0,1)$ as well?
Open Problem 6.2.
Let $(\mathrm{BP}(t))_{t\geq 0}$ be a CMJ branching process whose offspring and inter-birth times satisfy Assumption 4.1. If either $\sum_{i=1}^{\infty}X(i)<\infty$ with positive probability or $\widehat{\mu}(\lambda)=\infty$ for all $\lambda>0$, under what conditions do persistent $m$-hubs exist, or does persistence occur? Additionally, can one prove an equivalent result as in Theorems 4.2 or 4.4 without the assumption that $\widehat{\mu}(\lambda)<1$ for some $\lambda>0$?
Characterising the difference between the richest and oldest individual. In the ‘rich are old’ regime, we show that the difference between the birth-time of the richest and oldest individual in $(\mathrm{BP}(t))_{t\geq 0}$, that is, $I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}$, is a tight sequence of random variables with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$. It remains unclear if one can obtain a functional convergence of this random process.
Open Problem 6.3.
@@
Open Problem 6.4.
Suppose the sequences $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied. Assume that $\theta:=\lim_{t\to\infty}-\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)$ exists and recall $R$ from (2.14). If $\theta<R$ and Assumption (C-V), does persistence occur in the sense of (P)?
@@
Other modelling choices. In the (embedding of the) PAVD model the size of the offspring of an individual equals the number of children said individual has produced. Similarly, the rate at which an individual produces its next child or dies depends on the total number of children produced. Instead, one could count only alive children towards the offspring, or let the birth and death rates depend on the number of alive children. Either way of defining the model can be motivated from applications. However, the analysis for these alternative choices present more challenges, since now either an individual needs to produce $n$ children out of which $n-k$ should die to have an offspring of size $k$, or the rates depend not only on the parent individual but also on the lifetimes of its children. Deijfen concludes from simulations presented in [13] that the asymptotic behaviour of the degree distribution should not change significantly, and we expect this to be the case for other properties such as persistence as well. Still, it would be interesting to further investigate these models too.

### v2

@@ 5.3.2. There exists an old individual with large offspring
The next step is to show that there exists an ‘old’ individual that survives until time $\approx t$ and has a sufficiently large offspring. Combined with Proposition 5.7 and Lemma 5.8, this implies that the birth-time of the individual with the largest offspring (i.e. $I_{t}^{\mathrm{ct}}$) cannot be ‘too large’ compared to $O_{t}^{\mathrm{ct}}$. We make this precise in the following result.
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k\right)\geq 1-\varepsilon.$ (5.62)
@@
$\geq{}$ $\mathbb{P}_{\mathcal{S}}\big(\exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1}):D^{(v)}\geq k,S^{(v)}_{k}\leq t-A-u_{t}-\tfrac{1}{2}C_{1},S^{(v)}_{D^{(v)}+1}>t+A-u_{t}\big).$
@@
$1{}$ $-\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})|\leq m\right)$ (5.64)
@@
$\geq{}$ $\mathbb{P}\!\left(D\geq k\right)\mathbb{E}\left[\mathbbm{1}\{S_{k}\leq t-u_{t}-A-\tfrac{1}{2}C_{1}\}\exp\big(-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big)\right]$
$\times\exp\big(-(d^{*}+\eta)(2A+\tfrac{1}{2}C_{1})\big).$
@@
$\geq{}$ $\mathbb{E}\big[\mathbbm{1}\{S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]\}\exp\big(-(d^{*}+\eta)(t-u_{t}-A-\tfrac{1}{2}C_{1}-S_{k})\big)\big]$
$\geq{}$ $\mathbb{P}\!\left(S_{k}\in[t-u_{t}-A-C_{1},t-u_{t}-A-\tfrac{1}{2}C_{1}]\right)\mathrm{e}^{-\frac{1}{2}(d^{*}+\eta)C_{1}}.$
@@
$c_{\infty}(A,C_{1})\coloneq\frac{1}{2}\mathbb{P}\!\left(M_{\infty}\in[-A,-A+\tfrac{1}{2}C_{1}]\right),$ (5.69)
@@
$={}$ $d^{*}\Big[\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\Big]+\mathcal{K}_{\alpha}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big)+o(1).$
Using Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$, we can bound the right-hand side from above for all $t>t^{\prime\prime}$ by
@@
$\mathbb{P}_{\mathcal{S}}(\forall{}$ $s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k)$ (5.76)
@@
$={}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t}+\tfrac{1}{2}C_{1})|-|\mathcal{B}(0,u_{t})|\leq M^{-1}\exp\big(\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)\right)$
$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t}+\tfrac{1}{2}C_{1})|\exp\big(-\lambda^{*}(u_{t}+\tfrac{1}{2}C_{1})\big)\leq 2M^{-1}\right)$
@@
$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(\inf_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq 2M^{-1}\right)+\mathbb{P}\!\left(\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\geq M^{-1}\exp\big(\tfrac{\lambda^{*}}{2}C_{1}\big)\right).$
@@ 5.4. Persistence in the PAVD model
Fix $\varepsilon>0$ small, $A>0$ and $t^{\prime}>0$ large (according to Proposition 5.7, Lemma 5.8, and Lemma 5.9) and recall $k=k(C_{1})$ and $u_{t}$ from (5.50). Then, for any $t>t^{\prime}$:
@@
    We take $K>0$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ we have $O_{s}^{\mathrm{ct}}>u_{t}-K$.
@@
    We take $C_{1}>0$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ there exists $v\in\mathcal{B}(u_{t},u_{t}+\tfrac{1}{2}C_{1})\cap\mathcal{A}_{s}^{\mathrm{ct}}$ such that $\deg_{s}(v)\geq k$.
@@
    We take $C_{2}>\tfrac{1}{2}C_{1}$ sufficiently large so that, with probability at least $1-\varepsilon/3$, for all $s\in[t-A,t+A]$ and for all $v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}$ we have $\deg_{s}(v)<k$.
$(a)$, $(b)$, and $(c)$ follow from Proposition 5.7, Lemma 5.9, and Lemma 5.8, respectively. Then, $(b)$ and $(c)$ combined imply that $I_{s}^{\mathrm{ct}}\leq u_{t}+C_{2}$ for all $s\in[t-A,t+A]$, with probability at least $1-\tfrac{2}{3}\varepsilon$. Combined with $(a)$, we thus conclude that the result follows for $M=C_{2}+K$. ∎
We finally prove Theorem 2.6 by translating the results proved in this section for the CTPAVD branching process to the desired results for the discrete time PAVD model, leveraging Proposition 3.1.
@@
We let $F_{n,A}\coloneq\big[\frac{1}{\lambda^{*}}\log n-A,\frac{1}{\lambda^{*}}\log n+A\big]$ for ease of writing. Then,
@@
$\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(\forall s\in F_{n,A}:\Big|O_{s}^{\mathrm{ct}}-\frac{d^{*}}{\lambda^{*}(\lambda^{*}+d^{*})}\log n-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<M\right)$
@@
The second term on the right-hand side is at most $\varepsilon/3$ for all $n\geq n_{0}$ by (5.80). Using Proposition 5.7 with $t=\frac{1}{\lambda^{*}}\log n$ we can take $M=M(A,\varepsilon)$ and $n_{1}=n_{1}(A,\varepsilon)\geq n_{0}$ such that the first term on the right-hand side is at least $1-\varepsilon/3$ for all $n\geq n_{1}$, so that the entire lower bound is at least $1-\frac{2}{3}\varepsilon$. We then define
$\ell_{n}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}\log n+\frac{\lambda^{*}}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big),$ (5.82)
@@
$={}$ $\mathbb{P}_{\mathcal{S}}\left(O_{n}\in\big[\mathrm{e}^{-M^{\prime}+\ell_{n}},\mathrm{e}^{M^{\prime}+\ell_{n}}\big]\right)$
$\geq{}$ $\mathbb{P}_{\mathcal{S}}\left(O_{\tau_{n}}^{\mathrm{ct}}\in\Big[\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}},\frac{1}{\lambda^{*}}\ell_{n}+\frac{M^{\prime}}{2\lambda^{*}}\Big]\right)-\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$
$-\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}+\frac{M^{\prime}}{2\lambda^{*}}\Big)\geq\mathrm{e}^{M^{\prime}+\ell_{n}}\right).$
The first probability is at least $1-\frac{2}{3}\varepsilon$ by taking $M^{\prime}=2\lambda^{*}M$ and applying (5.81) with $M$ and $n$ sufficiently large. Then, since $N(t)\geq|\mathcal{B}((0,t))|$ for all $t\geq 0$ deterministically,
$\mathbb{P}_{\mathcal{S}}\left(N\Big(\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ $\leq\mathbb{P}_{\mathcal{S}}\left(\Big|\mathcal{B}\Big(0,\frac{1}{\lambda^{*}}\ell_{n}-\frac{M^{\prime}}{2\lambda^{*}}\Big)\Big|\leq\mathrm{e}^{-M^{\prime}+\ell_{n}}\right)$ (5.85)
@@
We continue by proving that $I_{n}/O_{n}$ is tight. We apply Proposition 5.10 and (5.80) to conclude that for any $\varepsilon>0$ there exist $n_{1}\geq n_{0}$ and $M=M(\varepsilon,A,n_{1})>0$ such that for all $n\geq n_{1}$,
@@
We can bound the first term on the right-hand side from below by $1-2\varepsilon/3$ for all $n\geq n_{1}$ using (5.86) and setting $M^{\prime}\coloneq C^{2}\mathrm{e}^{\lambda^{*}M}$. The second term can be bounded from above by using that $N(t)\leq 2|\mathcal{B}(0,t)|$ deterministically for all $t\geq 0$, so that
@@
$\geq{}$ $\mathbb{P}_{\mathcal{S}}\bigg(\forall s\in F_{n,A}:\Big|\varphi_{1}(\max_{v\in\mathcal{A}_{s}^{\mathrm{ct}}}\deg_{s}(v))-\frac{1}{\lambda^{*}+d^{*}}\log n+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}\big(r\big(\tfrac{1}{\lambda^{*}}\log n\big)\big)\Big|<K\bigg)$
@@
The final term is at most $\varepsilon/2$ for $A$ and $n$ sufficiently large by (5.80). The first probability on the right-hand side is also at least $1-\varepsilon/2$ by setting $C_{1}^{\prime}=K$ in Lemma 5.8 and $C_{1}=K$ in Lemma 5.9 and taking $K$ sufficiently large (with $t=\frac{1}{\lambda^{*}}\log n$ and $n$ sufficiently large). We thus arrive at the desired result, which concludes the proof. ∎
@@ 6. Discussion and Open Problems
Characterising the difference between the richest and oldest individual. In the ‘rich are old’ regime, we have showed that the difference between the birth-time of the richest and oldest individual in $(\mathrm{BP}(t))_{t\geq 0}$, that is, $I^{\mathrm{ct}}_{t}-O^{\mathrm{ct}}_{t}$, is a tight sequence of random variables with respect to the probability measure $\mathbb{P}_{\mathcal{S}}$. It remains unclear if more precise results can be proved.
@@
Open Problem 6.2.
Suppose the sequences $b$ and $d$ are such that Assumptions (N-E) and (Ma) are satisfied. Assume that $\theta\coloneq\lim_{t\to\infty}-\frac{1}{t}\log\mathbb{P}\!\left(L>t\right)$ exists and recall $R$ from (2.14). If $\theta<R$ and Assumption (C-V), does persistence occur in the sense of (P)?
@@
Other modelling choices. In the (embedding of the) PAVD model the offspring of an individual equals the number of children said individual has produced. Similarly, the rate at which an individual produces its next child or dies depends on the total number of children produced. Instead, one could let the birth and death rates depend on the number of alive children only. Either way of defining the model can be motivated from applications. However, analysing this alternative model presents more challenges, since now an individual needs to produce $n$ children out of which nkn-k should die to have an offspring of size $k$. The birth and death rates of an individual thus depend the lifetimes of its children. Deijfen concludes from simulations presented in [13] that the asymptotic behaviour of the degree distribution should not change significantly, and we expect this to be the case for other properties such as persistence as well. Still, it would be interesting to further investigate these models too.